OpenStudy (anonymous):

what is the square root of 200

OpenStudy (anonymous):

10

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\sqrt{2*100} = \sqrt{2}*\sqrt{100}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

not mension

OpenStudy (isaiah.feynman):

@tecateskid do you want it simplified?